4. An area of ​​2 m2 is covered by a 2 cm layer of ice. How much heat did the frozen water of the party release?
Answer #1

Heat released = m

mass of the ice m = density and volume

density of ice at 0 degree celsius = 916 kg/m^3

volume of the ice = 2*0.02 = 0.04 m^3  

mass of the water m = 916*0.04 = 36.64 kg

Heat releases   = m = 36.64 * 3.35*105 = 122.744 * 105 J = 12.27 * 106 J